OpenStudy (anonymous):

Simplify the equation if possible: 3x-5x+8=x-4+12

OpenStudy (anonymous):

Okay, so I combined my like terms and got: -2x+8=x+8 correct?

OpenStudy (anonymous):

So, then I got -3x+8=8

hartnn (hartnn):

yes correct so far

OpenStudy (anonymous):

-3x+8=8 -8 -8 ---------------- 0/-3x =0

OpenStudy (anonymous):

Would this be considered no solution, or just 0?

hartnn (hartnn):

yeah, you got -3x +8=8 -3x = 8-8 -3x=0 x = 0/-3 = 0 so 0 is a solution :)

OpenStudy (anonymous):

Okay, awesome thanks! :)

hartnn (hartnn):

welcome ^_^